# Prospects of constraining $f(T)$ gravity with the third–generation gravitational–wave detectors

### 03 Jan 2024 v1

@@ IV.1 BNS
∙$\bullet$ Injected redshift distribution −-- The redshift distribution of the BNS (NSBH) population follows Madau-Dickinson profile [61, 62], convolved with a time delay distribution $P\left(t_{d}\right)\propto 1/t_{d}$. Assuming a minimum time delay of 20 Myr, Iacovelli et al. [48] proposed an analytic equation (Eq. (A2) therein) to describe the distribution of redshift for these compact objects mergers. Therefore, we utilize it as the redshift distribution for our simulated GW events.
∙$\bullet$ Inclination angle $\theta_{\rm j}$-- The large uncertainty of $d_{\rm L}^{gw}$ arises from the degeneracy between $d_{\rm L}$ and $\iota$ in the GW events, which significantly impacts the constraints of cosmological models. Fortunately, the viewing angle inferred from SGRB is widely anticipated to be the same as the inclination angle of the BNS mergers. Therefore, we consider a specific case (on-axis) in which the line of sight is within the energetic core of the ejecta (i.e., $\theta_{\rm v}\leq\theta_{\rm j}$, $\theta_{\rm j}$ represents the evaluated jet opening angle and we assume $\theta_{\rm j}\sim 0.1\rm{rad}$). The uncertainty of the $\theta_{\rm v}$ (equally, the $\iota$) can be estimated to be within $\sim$ 0.1 rad. Another motivation for considering the on-axis case is that the afterglow emissions of off-axis SGRBs are highly suppressed until the blast waves driven by the ejecta have decelerated to a bulk Lorentz factor of $\Gamma=1/\theta_{\mathrm{v}}$ and reached the peak flux at that time. Thus, only the on-axis SGRBs have high detectability at high redshifts.
@@ IV.2 NSBH
Unlike BNS mergers, only a small proportion of NSBH mergers are expected to be accompanied with an EM counterpart. This phenomenon primarily occurs in cases where the stellar-mass black hole has a high spin during the merger event, allowing for the presence of residual baryon mass $M_{\mathrm{rem}}>0$ outside the black hole after the merger.
∙$\bullet$ BH mass $M_{BH}$ and spin $\chi_{BH}$-- Following Refs. [63, 44], we adopt a formula with two broken exponential decays and an exponential rise to describe the BH mass distribution, i.e.,
@@
where $M_{\rm min}=5M_{\odot}$ is the minimum BH mass based on the rapid core-collapse supernova model proposed by Ref. [64]. The best-fit values for the other parameters are fixed $a_{1}=1.04\times 10^{11}$, $b_{1}=2.1489$, $a_{2}=799.1$, $b_{2}=0.2904$, $a_{3}=0.002845$, and $b_{3}=1.686$ [44].
∙$\bullet$ NS mass $M_{NS}$ and spin $\chi_{NS}$-- As shown in population synthesis studies [63], the masses of NSs in NSBH systems may be larger than those in BNS systems. The median value of NS masses in NSBH systems is around $1.6M_{\odot}$. Therefore, we assume a normal distribution $M_{\mathrm{NS}}/M_{\odot}\sim\mathcal{N}\left(1.6,0.11^{2}\right)$ with the AP4 NS equation of state (EoS). Regarding the NS spin, we consider a low-spin case since the largest spin of a NS measured in a binary system is 0.02 [65, 66], and we assume $\chi_{\mathrm{NS}}\sim\text{Uniform}\left(0,0.05\right)$.
∙$\bullet$ Remnant baryon mass −--In order to determine whether an EM counterpart is present in a NSBH merger, it is crucial to estimate the remnant baryon mass $\left(M_{\mathrm{rem}}\right)$ outside the black hole after the merger. A positive value of $M_{\mathrm{rem}}$ indicates the formation of a disk that can generate EM counterparts, such as kilonovae, gamma-ray bursts (GRBs), and afterglows [67]. On the other hand, if $M_{\mathrm{rem}}\leqslant 0$, it suggests that the NS plunges directly into the black hole without leaving any observable signatures except for the GW.
In the simulations, we focus on the high-spin case for the black hole and assume a normal distribution for the dimensionless spin parameter $\chi_{\mathrm{BH}}$, given by $\chi_{\mathrm{BH}}\sim\mathcal{N}\left(0.8,0.15^{2}\right)$. This choice is consistent with observations in black hole X-ray binary systems [68, 69]. The remnant disk mass is estimated using the method described in Ref. [60]
@@ IV.3 Detectability
where $S_{n}(f)$ is the one-sided power spectral density of the GW detector. The considered frequency range is from $20$ Hz to $2048$ Hz. The net SNR of the network consisting of ET+2CE detectors, as well as the root sum squared of the SNR of all detectors, are calculated using the Bilby [72] software package. If the net SNR exceeds the threshold value of $\rho_{*}=12$, the GW signal is assumed to be reliably detected and can be utilized for further analysis. The redshift distributions of our simulated BNS and NSBH events are presented in Fig. 3.
@@
Hubble constant $H_{0}\left[\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}\right]$ $68.6$ Uniform$(48.6,88.6)$
@@
Parameter of $f(T)$ gravity $\beta$ 00 Uniform(−0.5,0.5)0.50.5(-0.5,0.5)( - 0.5 , 0.5 ) [73, 14]a
Absolute magnitude $M$$-19.253$ Uniform(−$(-18.253,-20.253)$ [74]a
@@
IV.4 Generation of standard sirens
Figure 4: The relation between GW luminosity distance and redshift. The solid black line represents the $\Lambda$CDM model, while the dotted and dashdot black lines represent the $f(T)$ model with $\beta=-0.3$ and $\beta=0.3$, respectively. The red and blue error bars represent the estimated $d_{L}^{gw}$ with 68% credible level (C.L.) for simulated BNS and NSBH events, respectively. The number of these two types standard sirens is equal to 100.
The corresponding luminosity distances ($d_{L}\left(z_{i}\right)$) of the simulated BNS and NSBH events are generated based on the $\Lambda\mathrm{CDM}$ model. The values of the cosmological parameters $H_{0}$ and $\Omega_{m}^{(0)}$ are set as 68.6 and 0.30, respectively, as constrained by SN Ia and BAO data.
$L$ =∏$=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi}\sigma_{i}\left(z_{i}\right)}\exp\left(-\frac{1}{2}\chi_{\mathrm{GW}}^{2}\right),$
where $d_{L}^{gw}$ is the GW “luminosity distance” in $f(T)$ gravity, as defined by Eq. (11). The subscript $i$ denotes the $i$-th event in the simulation catalog.
@@ V RESULTS AND DISCUSSION
$\mathrm{SNIa}+\mathrm{BAO}(\Lambda\rm CDM)$ $68.6_{-0.8}^{+0.7}$ $0.300_{-0.020}^{+0.024}$-- −$-19.392_{-0.021}^{+0.019}$ $146.85_{-0.04}^{+1.32}$
@@
$\mathrm{BNS(100)}$ $68.5_{-0.7}^{+0.6}$ $0.231_{-0.064}^{+0.165}$ $0.17_{-0.41}^{+0.16}$-- −--
$\mathrm{BNS(1000)}$ $69.0_{-0.3}^{+0.4}$ $0.273_{-0.079}^{+0.183}$ $0.22_{-0.38}^{+0.15}$-- −--
$\mathrm{NSBH(100)}$ $68.7_{-0.7}^{+0.6}$ $0.376_{-0.089}^{+0.124}$ $0.25_{-0.33}^{+0.28}$-- −--
$\mathrm{NSBH(1000)}$ $68.8_{-0.3}^{+0.4}$ $0.303_{-0.038}^{+0.071}$ $0.06_{-0.12}^{+0.14}$-- −--
@@
Here, we present the results of parameter estimations obtained using Monte Carlo Markov Chains (MCMCs) and discuss the potential of constraining $f(T)$ gravity using the simulated GW events. To obtain the posterior distributions for cosmological parameters, we employ emcee [79] as the sampler. Additionally, we utilize the getdist [80] code to facilitate the visualization of the chains and generate plots from the final results. All of the prior distributions of the cosmological parameters are represented in Table 2. Within the $\Lambda\mathrm{CDM}$ framework, fiducial values for these parameters are obtained from analyzing current observations, including SN Ia, and BAO, with best fitting values.
Figure 5: Posterior distributions of the cosmological parameters listed in Table 3 utilizing GW events only. Using 100 (1000) BNS and NSBH events, the results of the Power-low model of $f(T)$ gravity are marked in red, green, orange, and blue, respectively. Red stars represent the fiducial values $\left(H_{0},\Omega_{m}^{(0)},\beta\right)=(68.6,0.30,0)$. The contours are at the 68%, 95% C.L.
As shown in Fig. 4, the discrepancies between $d_{\rm L}^{gw}$ and $d_{\rm L}^{\mathrm{em}}$ caused by parameters changing under $f(T)$ gravity become remarkable at higher redshifts. Consequently, the NSBH samples are anticipated to offer more stringent constraints on $f(T)$ gravity compared to the BNS samples, as illustrated in Fig. 5. Moreover, there is a substantial degeneracy between the parameters of $\Omega_{m}^{(0)}$ and $\beta$ for analysis of both binary systems. This degeneracy arises from the combined term of $\beta$ and $\Omega_{m}^{(0)}$ in Eq. (6)and Eq. (12). The former equation plays a significant role in the analysis of SN Ia and BAO data and the latter one dominates the analysis of GW, generating different degeneracy in Fig. 5 and Fig. 6.
Figure 6: Posterior distributions of the parameters listed in Table 3 utilizing GW events, SN Ia, and BAO observations. The results for the Power-law model of $f(T)$ gravity, obtained using different combinations of data, are represented by different colors. Red stars represent the fiducial values $\left(H_{0},\Omega_{m}^{(0)},\beta\right)=(68.6,0.30,0)$. The contours are at the 68%, 95% C.L.
The results obtained from combining GW simulations, SN Ia, and BAO observations are shown in Table 3 and Fig. 6. Compared to the analysis of SN Ia and BAO observations alone, which yielded $\Omega_{m}^{(0)}=0.331_{-0.061}^{+0.034}$ and $\beta=-0.35_{-0.10}^{+0.49}$ at 68% credible level (C.L.), the combined analysis breaks the degeneracy between these two parameters and provides more precise estimates for them. Specifically, the combined analysis from SN Ia + BAO + NSBH (1000) yields $\Omega_{m}^{(0)}=0.300_{-0.011}^{+0.011}$ and $\beta=0.005_{-0.030}^{+0.028}$ at 68% C.L.
@@ VI Conclusion
The developing field of multi-messenger astronomy, facilitated by GWs, provides a novel avenue for investigating gravity and cosmology. In this work, we focus on utilizing the simulated BNS and NSBH events to prospectively constrain $f(T)$ gravity with the third-generation gravitational-wave detectors.
Based on current observations, we conducted a comprehensive simulation of the BNS and NSBH mergers, taking into account realistic selection criteria to ensure that these simulated events are always accompanied with electromagnetic counterparts with certain redshifts. Moreover, we restrict the inclination angle to a small range (∼$\sim 0.1\mathrm{rad}$) to enhance the realism and precision of uncertainty estimation in the GW luminosity distance. The expected deviations in the ratio $d_{L}^{gw}/d_{L}^{em}$ caused by $f(T)$ gravity are anticipated to exceed 15% at high redshifts ($z\sim 5$). Therefore, we can compare the modified $d_{\rm L}^{gw}$ with $d_{\rm L}^{em}$ to constrain $f(T)$ gravity and other cosmological parameters.
By analyzing SN Ia and BAO data, we established the baseline of simulated GW events in the frame of $\Lambda\mathrm{CDM}$ and constrained the parameter $\beta=-0.35_{-0.10}^{+0.49}$ at 68% C.L. in the Power-law model under $f(T)$ gravity. Moreover, we incorporated simulated GW events into our analysis, combined with SN Ia and BAO observations, and discovered that this combination effectively breaks the degeneracy between $\Omega_{m}^{(0)}$ and $\beta$, leading to the tightest constraints that $\beta=0.005_{-0.030}^{+0.028}$ at 68% C.L. This improved precision shows an order of magnitude greater than the analyzing result based on SN Ia and BAO data, which enables us to determine the deviation between modified gravity and GR. In addition, the GW simulated events provide $H_{0}$ with uncertainties ranging from 1.0% to 1.9%, indicating a promising avenue to explore the $H_{0}$ tension under the $f(T)$ gravity framework.
@@
Acknowledgments.
— We are grateful to Wen Zhao, ShaoPeng Tang, MingZhe Han, Rui Niu, Zhao Li, Jin Qiao, Xin Ren, and BinBin Zhang for helpful discussions. This work is supported in part by the National Natural Science Foundation of China (No. 12233011 and 11921003).

### 30 Jan 2024 v2

@@ IV.1 BNSs
(iii) Injected redshift distribution: The redshift distribution of the BNS (NSBH) population follows Madau-Dickinson profile [61, 62], convolved with a time delay distribution $P\left(t_{d}\right)\propto 1/t_{d}$. Assuming a minimum time delay of 20 Myr, Iacovelli $et$ $al$. [48] proposed an analytic equation [Eq. (A2) therein] to describe the distribution of redshift for these compact objects mergers. Therefore, we utilize it as the redshift distribution for our simulated GW events.
(iv) Inclination angle $\theta_{\rm j}$: The large uncertainty of $d_{\rm L}^{gw}$ arises from the degeneracy between $d_{\rm L}$ and $\iota$ in the GW events, which significantly impacts the constraints of cosmological models. Fortunately, the viewing angle inferred from SGRBs is widely anticipated to be the same as the inclination angle of the BNS mergers. Therefore, we consider a specific case (on axis) in which the line of sight is within the energetic core of the ejecta (i.e., $\theta_{\rm v}\leq\theta_{\rm j}$, $\theta_{\rm j}$ represents the evaluated jet opening angle and we assume $\theta_{\rm j}\sim 0.1$ $\rm{rad}$). The uncertainty of the $\theta_{\rm v}$ (equally, the $\iota$) can be estimated to be within $\sim$ 0.1 rad. Another motivation for considering the on axis case is that the afterglow emissions of off axis SGRBs are highly suppressed until the blast waves driven by the ejecta have decelerated to a bulk Lorentz factor of $\Gamma=1/\theta_{\mathrm{v}}$ and reached the peak flux at that time. Thus, only the on axis SGRBs have high detectability at high redshifts.
@@ IV.2 NSBH
Unlike BNS mergers, only a small proportion of NSBH mergers are expected to be accompanied with an EM counterpart. This phenomenon primarily occurs in cases where the stellar-mass black hole has a high spin during the merger event, allowing for the presence of residual baryon mass $M_{\mathrm{rem}}>0$ outside the black hole (BH) after the merger.
(i) BH mass $M_{BH}$ and spin $\chi_{BH}$: Following Refs. [63, 44], we adopt a formula with two broken exponential decays and an exponential rise to describe the BH mass distribution, i.e.,
@@
where $M_{\rm min}=5\mathrm{M}_{\odot}$ is the minimum BH mass based on the rapid core-collapse supernova model proposed by Ref. [64]. The best-fit values for the other parameters are fixed $a_{1}=1.04\times 10^{11}$, $b_{1}=2.1489$, $a_{2}=799.1$, $b_{2}=0.2904$, $a_{3}=0.002845$, and $b_{3}=1.686$ [44].
(ii) NS mass $M_{NS}$ and spin $\chi_{NS}$: As shown in population synthesis studies [63], the masses of NSs in NSBH systems may be larger than those in BNS systems. The median value of NS masses in NSBH systems is around $1.6\mathrm{M}_{\odot}$. Therefore, we assume a normal distribution $M_{\mathrm{NS}}/\mathrm{M}_{\odot}\sim\mathcal{N}\left(1.6,0.11^{2}\right)$ with the AP4 NS equation of state (EoS). Regarding the NS spin, we consider a low-spin case since the largest spin of a NS measured in a binary system is 0.02 [65, 66], and we assume $\chi_{\mathrm{NS}}\sim\text{uniform}\left(0,0.05\right)$.
(iii) Remnant baryon mass: In order to determine whether an EM counterpart is present in a NSBH merger, it is crucial to estimate the remnant baryon mass $\left(M_{\mathrm{rem}}\right)$ outside the black hole after the merger. A positive value of $M_{\mathrm{rem}}$ indicates the formation of a disk that can generate EM counterparts, such as kilonovae, GRBs, and afterglows [67]. On the other hand, if $M_{\mathrm{rem}}\leqslant 0$, it suggests that the NS plunges directly into the black hole without leaving any observable signatures except for the GW.
In the simulations, we focus on the high-spin case for the black hole and assume a normal distribution for the dimensionless spin parameter $\chi_{\mathrm{BH}}$, given by $\chi_{\mathrm{BH}}\sim\mathcal{N}\left(0.8,0.15^{2}\right)$. This choice is consistent with observations in black hole x-ray binary systems [68, 69]. The remnant disk mass is estimated using the method described in Ref. [60],
@@ IV.3 Detectability
where $S_{n}(f)$ is the one-sided power spectral density of the GW detector. The considered frequency range is from $20$ to $2048$ Hz. The net SNR of the network consisting of ET+2CE detectors, as well as the root sum squared of the SNR of all detectors, is calculated using the Bilby [72] software package. If the net SNR exceeds the threshold value of $\rho_{*}=12$, the GW signal is assumed to be reliably detected and can be utilized for further analysis. The redshift distributions of our simulated BNS and NSBH events are presented in Fig. 3.
IV.4 Generation of standard sirens
Figure 4: The relation between GW luminosity distance and redshift. The solid black line represents the $\Lambda$CDM model, while the dotted and dash-dotted black lines represent the $f(T)$ model with $\beta=-0.3$ and $\beta=0.3$, respectively. The red and blue error bars represent the estimated $d_{L}^{gw}$ with 68% credible level (CL.) for simulated BNS and NSBH events, respectively. The number of these two types of standard sirens is equal to 100.
The corresponding luminosity distances [$d_{L}\left(z_{i}\right)$] of the simulated BNS and NSBH events are generated based on the $\Lambda\mathrm{CDM}$ model. The values of the cosmological parameters $H_{0}$ and $\Omega_{m}^{(0)}$ are set as 68.6 and 0.30, respectively, as constrained by SN Ia and BAO data.
$L$ $=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi}\sigma_{i}\left(z_{i}\right)}\exp\left(-\frac{1}{2}\chi_{\mathrm{GW}}^{2}\right),$
where $d_{L}^{gw}$ is the GW “luminosity distance” in $f(T)$ gravity, as defined by Eq. (11). The subscript $i$ denotes the $i$th event in the simulation catalog.
@@ IV.4 Generation of standard sirens
Hubble constant $H_{0}\left[\mathrm{km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}\right]$ $68.6$ Uniform$(48.6,88.6)$
@@
Parameter of $f(T)$ gravity $\beta$ 00 Uniform(0.5,0.5)0.50.5(-0.5,0.5)( - 0.5 , 0.5 ) [76, 14]a
Absolute magnitude $M$ $-19.253$ Uniform$(-18.253,-20.253)$ [77]a
@@ V RESULTS AND DISCUSSION
Here, we present the results of parameter estimations obtained using Monte Carlo Markov chains and discuss the potential of constraining $f(T)$ gravity using the simulated GW events. To obtain the posterior distributions for cosmological parameters, we employ emcee [79] as the sampler. Additionally, we utilize the getdist [80] code to facilitate the visualization of the chains and generate plots from the final results. All of the prior distributions of the cosmological parameters are represented in Table 2. Within the $\Lambda\mathrm{CDM}$ framework, fiducial values for these parameters are obtained from analyzing current observations, including SN Ia and BAOs, with best fitting values.
As shown in Fig. 4, the discrepancies between $d_{\rm L}^{gw}$ and $d_{\rm L}^{\mathrm{em}}$ caused by parameters changing under $f(T)$ gravity become remarkable at higher redshifts. Consequently, the NSBH samples are anticipated to offer more stringent constraints on $f(T)$ gravity compared to the BNS samples, as illustrated in Fig. 5. Moreover, there is a substantial degeneracy between the parameters of $\Omega_{m}^{(0)}$ and $\beta$ for analysis of both binary systems. This degeneracy arises from the combined term of $\beta$ and $\Omega_{m}^{(0)}$ in Eqs. (6) and (12). The former equation plays a significant role in the analysis of SN Ia and BAO data and the latter one dominates the analysis of GWs, generating different degeneracy in Figs. 5 and 6.
@@
$\mathrm{SNIa}+\mathrm{BAO}(\Lambda\rm CDM)$ $68.6_{-0.8}^{+0.7}$ $0.300_{-0.020}^{+0.024}$- $-19.392_{-0.021}^{+0.019}$ $146.85_{-0.04}^{+1.32}$
@@
$\mathrm{BNS(100)}$ $68.5_{-0.7}^{+0.6}$ $0.231_{-0.064}^{+0.165}$ $0.17_{-0.41}^{+0.16}$- --
$\mathrm{BNS(1000)}$ $69.0_{-0.3}^{+0.4}$ $0.273_{-0.079}^{+0.183}$ $0.22_{-0.38}^{+0.15}$- --
$\mathrm{NSBH(100)}$ $68.7_{-0.7}^{+0.6}$ $0.376_{-0.089}^{+0.124}$ $0.25_{-0.33}^{+0.28}$- --
$\mathrm{NSBH(1000)}$ $68.8_{-0.3}^{+0.4}$ $0.303_{-0.038}^{+0.071}$ $0.06_{-0.12}^{+0.14}$- --
@@
Figure 5: Posterior distributions of the cosmological parameters listed in Table 3 utilizing GW events only. Using 100 (1000) BNS and NSBH events, the results of the power-low model of $f(T)$ gravity are marked in red, green, orange, and blue, respectively. Red stars represent the fiducial values $\left(H_{0},\Omega_{m}^{(0)},\beta\right)=(68.6,0.30,0)$. The contours are at the 68% and 95% CL.
The results obtained from combining GW simulations, SN Ia, and BAO observations are shown in Table 3 and Fig. 6. Compared to the analysis of SN Ia and BAO observations alone, which yielded $\Omega_{m}^{(0)}=0.331_{-0.061}^{+0.034}$ and $\beta=-0.35_{-0.10}^{+0.49}$ at 68% CL., the combined analysis breaks the degeneracy between these two parameters and provides more precise estimates for them. Specifically, the combined analysis from SN Ia + BAO + NSBH (1000) yields $\Omega_{m}^{(0)}=0.300_{-0.011}^{+0.011}$ and $\beta=0.005_{-0.030}^{+0.028}$ at 68% CL.
@@
Figure 6: Posterior distributions of the parameters listed in Table 3 utilizing GW events, SN Ia, and BAO observations. The results for the power-law model of $f(T)$ gravity, obtained using different combinations of data, are represented by different colors. Red stars represent the fiducial values $\left(H_{0},\Omega_{m}^{(0)},\beta\right)=(68.6,0.30,0)$. The contours are at the 68% and 95% CL.
@@ VI Conclusion
The developing field of multimessenger astronomy, facilitated by GWs, provides a novel avenue for investigating gravity and cosmology. In this work, we focus on utilizing the simulated BNS and NSBH events to prospectively constrain $f(T)$ gravity with the third-generation gravitational-wave detectors.
Based on current observations, we conducted a comprehensive simulation of the BNS and NSBH mergers, taking into account realistic selection criteria to ensure that these simulated events are always accompanied with electromagnetic counterparts with certain redshifts. Moreover, we restrict the inclination angle to a small range ($\sim 0.1$ $\mathrm{rad}$) to enhance the realism and precision of uncertainty estimation in the GW luminosity distance. The expected deviations in the ratio $d_{L}^{gw}/d_{L}^{em}$ caused by $f(T)$ gravity are anticipated to exceed 15% at high redshifts ($z\sim 5$). Therefore, we can compare the modified $d_{\rm L}^{gw}$ with $d_{\rm L}^{em}$ to constrain $f(T)$ gravity and other cosmological parameters.
By analyzing SN Ia and BAO data, we established the baseline of simulated GW events in the frame of $\Lambda\mathrm{CDM}$ and constrained the parameter $\beta=-0.35_{-0.10}^{+0.49}$ at 68% CL. in the power-law model under $f(T)$ gravity. Moreover, we incorporated simulated GW events into our analysis, combined with SN Ia and BAO observations, and discovered that this combination effectively breaks the degeneracy between $\Omega_{m}^{(0)}$ and $\beta$, leading to the tightest constraints that $\beta=0.005_{-0.030}^{+0.028}$ at 68% CL. This enhanced precision shows an order of magnitude improvement over the analyzing result based on SN Ia and BAO data, which enables us to determine the deviation between modified gravity and GR. In addition, the GW simulated events provide $H_{0}$ with uncertainties ranging from 1.0% to 1.9%, indicating a promising avenue to explore the $H_{0}$ tension under the $f(T)$ gravity framework.
@@
Acknowledgements.
We are grateful to Wen Zhao, ShaoPeng Tang, MingZhe Han, Rui Niu, Zhao Li, Jin Qiao, Xin Ren, and BinBin Zhang for helpful discussions. This work is supported in part by the National Natural Science Foundation of China (No. 12233011 and 11921003).